OpenStudy (diyadiya):

Find all points of discontinuity of f, where f is defined by :

OpenStudy (mr.math):

Someone's studying!

OpenStudy (diyadiya):

\[\large \binom{\begin{cases} |x|+3& \text{ if } x\leq -3 \\ -2x & \text{ if } -3<x<3 \\ 6x+2 & \text{ if } x\geq 3 \end{cases}}{}\]

OpenStudy (mr.math):

Let's start with x=-3, okay?

OpenStudy (mr.math):

Can you tell me what's the limit as x approaches \(x^-\)?

OpenStudy (diyadiya):

ok

OpenStudy (diyadiya):

isn't it \[3^-\]

OpenStudy (mr.math):

Yeah, I meant \(3^-\), what's the lmit?

OpenStudy (diyadiya):

|x|+3 so -x+3 & x+3 right?

OpenStudy (mr.math):

Note that \(|x|+3=-x+3\) for \(x\le -3\).

OpenStudy (mr.math):

She's a girl?

OpenStudy (mr.math):

:P

OpenStudy (mr.math):

Anyways, what did you find for the limit?

OpenStudy (anonymous):

I gave her a medal.

OpenStudy (diyadiya):

why is it just -x+3 ? what about x+3 ? @Fool I know!!

OpenStudy (mr.math):

Because x<=-3, right?

OpenStudy (anonymous):

because of the constraints ..

OpenStudy (anonymous):

diya you are not thinking :/

OpenStudy (diyadiya):

oh ok

OpenStudy (anonymous):

That's not good thing ... Limit is a useful thing to understand ...

OpenStudy (mr.math):

I will help you with this one. FOCUS! \[\lim_{x\to -3^-}f(x)=\lim_{x\to-3^-}|x|+3=\lim_{x\to-3^-}-x+3-=-(-3)+3=6.\]

OpenStudy (mr.math):

Can you find the limit as x approaches \(-3^+\)?

OpenStudy (diyadiya):

wait wait

OpenStudy (mr.math):

\[\lim_{x\to -3^+}f(x)=\lim_{x\to -3^+}-2x=...\]

OpenStudy (diyadiya):

-2(-3)=6

OpenStudy (mr.math):

Good! Are they the same?

OpenStudy (mr.math):

I mean the left and right limit as x ->3?

OpenStudy (diyadiya):

lol Yes

OpenStudy (mr.math):

What do you conclude then?

OpenStudy (diyadiya):

its continuos at -3

OpenStudy (mr.math):

What about f(-3)?

OpenStudy (diyadiya):

f(-3)=-(-3)+3=6

OpenStudy (mr.math):

Now, you can say that f(x) is continuous at x=-3.

OpenStudy (mr.math):

Try to do the same with x=3, and write down what you get.

OpenStudy (diyadiya):

ok

OpenStudy (anonymous):

diya just when right hand limit and let hand limits are equal, we can only say that that the limit exist at that point.

OpenStudy (mr.math):

Find left limit, right limit and the value of the function at that point. If they are all the same then it's continuous at that point.

OpenStudy (diyadiya):

yeah i know that :)

OpenStudy (mr.math):

Because you're smart! :P

OpenStudy (mr.math):

Why does Diya have two medals and I only have one?! ::P

OpenStudy (diyadiya):

ok wait i'll do the rest

OpenStudy (diyadiya):

beacuse diya gives medals in the end!

OpenStudy (mr.math):

lol

OpenStudy (diyadiya):

\[\lim_{x \rightarrow 3^+}f(x)=\lim_{x \rightarrow 3^+}(6x+2)=20\] \[\lim_{x \rightarrow 3^-}f(x)=\lim_{x \rightarrow 3^-}-2x=-6\] \[f(3)=6(3)+2=20\] \[LHL \neq RHL\] So its discontionous at 3 Right or wrong ?

OpenStudy (mr.math):

Awesome!

OpenStudy (mr.math):

You deserve a medal NOW!

OpenStudy (diyadiya):

Lol okay ofcourse You too!!

OpenStudy (diyadiya):

\[\huge THANK YOU \]

OpenStudy (diyadiya):

to both of you!

OpenStudy (anonymous):

analyze how this incident will affect the over all automobile industry? diya can u please translate it in urdu >

OpenStudy (anonymous):

nobody urdu here mam...

OpenStudy (diyadiya):

i dont know urdu sorry

OpenStudy (anonymous):

oh:( jhooti